# Hesitation and Tolerance in Recommender Systems

### v1

@@ 2.3. User Retention Optimization
User retention is typically defined as the ratio of users who revisit a system after their initial interaction (Cai et al., 2023). It has become an increasingly critical metric in the evaluation and optimization of recommender systems (Wu et al., 2017; Zou et al., 2019; Cai et al., 2023; Xue et al., 2023b; Silveira et al., 2017; Zhang et al., 2021b; Ding et al., 2023a; Gomez-Uribe and Hunt, 2016). Unlike immediate feedback mechanisms, such as point-wise (He et al., 2021; Zamani and Croft, 2020) and list-wise (Cui et al., 2021; Zheng et al., 2020) signals, which reflect user responses after a single interaction, retention captures long-term feedback resulting from multiple interactions over time (Cai et al., 2023).
@@
Due to the nature of this delayed feedback, reinforcement learning (RL) methods are particularly well-suited to optimize user retention by addressing long-term rewards and user engagement (Zhang et al., 2022; Xue et al., 2023a; Zou et al., 2019). One of the representative works is RLUR (Cai et al., 2023), designed to address challenges such as uncertainty, bias, and delayed feedback, thereby improving retention by focusing on long-term positive outcomes. In addition, some studies have explored optimizing retention through RL by modeling mid- and long-term user behaviors (Wu et al., 2017; Zou et al., 2019; Xue et al., 2023b), investigating interpretable aspects of retention (Ding et al., 2023b), or integrating exploratory behaviors with reinforcement learning to enhance user experiences (Chen et al., 2021). Recent efforts (Kohavi and Chen, 2024) also analyze the impact of false positives in online A/B tests for recommender systems, where the mined false positives can be regarded as the retention signal.
3. User Behavior understanding through Surveys
In recommender systems, user preferences are often inferred from interactions with items, enabling refinement of the recommendation model. However, this inference process can introduce undesirable biases and errors. For instance, clicking on an item does not always indicate genuine interest; it may simply reflect a desire for more information. When users explore item details only to realize a lack of interest, interpreting the initial click as a preference signal leads to inaccurate assumptions. Consequently, such actions produce misleading feedback, potentially increasing uncertainty in future recommendations.
In addition, the decision-making process among users, based on their stated preferences of “interested” or “disinterested”, involves a series of complex and varied emotional processes. Before solidifying into definitive attitudes of interest and disinterest, individuals may incur certain costs, including time and potentially financial resources. For example, consider a scenario where a platform recommends a video that gradually loses the user’s interest until they eventually dismiss it. This scenario contrasts with a situation where the user immediately expresses disinterest and swipes away the recommended video. Determining which scenario yields a superior user experience is challenging. While the former scenario incurs a greater time cost to the user, the latter facilitates rapid disengagement based on disinterest. Consequently, there is a need to emphasize the investigation of emotional dynamics during user decision-making. This involves exploring potential novel emotional factors and their consequential influence on recommender systems.
3.1. User Survey Hypothesis
Based on the above discussion, we propose the following 4 hypotheses for our research objectives, and we verify them through user survey.
  • In the utilization of recommender systems, there exists an intermediary decision-making state, described as “hesitation”. Hesitation occurs before users can decide whether they are interested or not, and takes more costs than immediate decision.
  • Hesitation constitutes a state of the process, yet the incurred cost of hesitation may affect a user’s level of interest or disinterest in an item.
  • Hesitation is associated with increased decision-making costs. When users, following a period of hesitation, conclude that they are disinterested, this transition often gives rise to an emotion termed as “tolerance”. This emotional state is characterized by a more negative user experience compared to immediate disinterest.
  • The accumulation of tolerance results in a diminution of user trust, consequently diminishing interest in the recommender system and leading to a decrease in user activity and ultimately resulting in user attrition.
Figure 1. The potential relationship between users’ trust in the platform and their interest in its recommended items.
These hypotheses are expected to shed light on the intricate process of user emotional responses during interactions with recommender systems, potentially uncovering subtle factors that could impact the systems’ efficiency in retaining users. An illustration of our hypotheses is provided in Figure 1. By comprehending and addressing these intermediate states, particularly hesitation, and tolerance, there is potential to optimize recommender systems to better align with user requirements, and enhance user trust and retention rates within the platform.
3.2. Survey Design Principles
We conduct user survey to test the above hypotheses. It is crucial to meticulously design the survey methodology (Krosnick, 2018; Schwarz, 1999). Notably, a well-designed survey should strategically avoid many direct questions about hesitation, as they often lead to inaccurate and superficial responses, and incorrect survey results may occur due to the inclusion of erroneous options. Also, although we introduce the concept of “tolerance” in this study, we cannot use this term directly in surveys. Instead, we aim to capture tolerant behavior by describing specific scenarios and combinations of user actions, such as “assuming interest but realizing not, after further exploration” to examine its impact. Further, to avoid potential biases, we conduct two separate user surveys. Each survey is divided into five distinct sections of questions, with each section designed to sequentially capture users’ thought processes, ensuring the reliability and validity of the results. Additionally, the composition of each question set aligns with the core idea behind the questionnaire design.
In the first user survey, we collect common user behaviors that express interest and disinterest, as well as behaviors where users initially think they are interested but later realize they are not. We also assess users’ attitudes by asking them to judge the extent to which they agreed with each of our predefined answers, thereby verifying the reasonableness of our designed options. At the same time, we also investigate whether users have experienced hesitation and explore the reasons they believe contributed to it. Additionally, we outline a series of scenarios where users realize they are actually not interested after hesitating. Users are asked to evaluate these scenarios based on their own experiences and choose the ones that have impacted their use of recommender systems.
In the second user survey, we design the first three sections of questions with the same objectives as the first survey to cross-verify the validity of the conclusions drawn from the two user surveys. We then introduce two new sections of questions that focus on the impact of hesitation and tolerance, such as whether hesitation affects the level of user interest, whether increase in tolerance behaviors decreases user engagement, and ultimately leads to user attrition.
Note that the two user surveys are conducted independently at different time periods (detailed in Section 3.5). We do not intentionally ask the same group of users who completed the first survey to answer the second one. In each survey, we include a set of basic questions to understand the demographic profile of the participants, such as age, occupation, and the highest level of education. This demographic information helps us assess the distribution of both surveys, determine whether the traffic on the survey platform is healthy, and evaluate the validity of the conclusions drawn.
3.3. Design of the First User Survey
The first user survey is structured into five sections, each serving a distinct purpose as described below. The survey questions are listed in Appendix A.1.
User profile. As a required option in the survey questionnaire, we collect user’s basic information such as age, education, and occupation in this section to get a general understanding of our respondents and to evaluate the quality of the survey platform where the questionnaire is posted.
Context Introduction. This section initiates by posing elementary inquiries aimed at engaging the participants within the context of utilizing recommender systems. Participants are prompted to reflect upon their experiences with recommender systems, including aspects they found favorable and areas they identified as lacking. These inquiries are instrumental in elucidating the perceived benefits of recommender systems to users. They constitute a foundational element for the subsequent survey segments as values.
User Behavior Survey. We select two scenarios, online shopping and content browsing (in the context of short-video sharing), to collect users’ interactive behaviors when they are interested, not interested, or initially think they are interested but later realize they are not, in the recommended results. Additionally, we consider the potential impact of incidental behavior and trending topics on the recommendation results, aiming to verify whether these factors influence the recommendation outcomes and whether they harm the user experience.
Description and Assessment of the “Hesitation” State. We design questions to inquire whether users have experienced a state of hesitation regarding recommendation results, and what might be the reasons behind. Through these questions, we aim to help users recall their past experiences on recommender systems, thereby concretizing the psychological behavior of “hesitation”. The answers facilitate our subsequent understanding of user decisions and corresponding psychological states following the “hesitation” phase.
Recognition of Behavioral Combinations. We create a series of hypothetical scenarios, primarily within the contexts of online shopping and content browsing, in which users encounter recommendation results that they initially believe to be of their interest but later realize are not. We ask users to identify whether these behavior combinations reflect their specific psychological states based on their own experiences. The options are designed to be simple and straightforward, with the aim of obtaining reliable results.
3.4. Design of the Second User Survey
In the second survey, the first three sections include the same questions as those in the first survey to validate responses from different user groups. Recall that the two surveys are designed to be conducted separately at different time periods. The remaining two sections of the second survey focus more on the impact of hesitation. Specifically, we have the following two sections of new questions in the second user survey (see Appendix A.2 for the list of questions).
Further Investigation into Hesitation. In this section, the questions are designed to ask whether users generally prefer items they have thought about carefully (i.e., experienced hesitation before realizing they like them), or if they trust their first instinct and prefer items that do not cause hesitation, we aim to test the fourth hypothesis in Section 3.1, which is the inverse relationship between hesitation behavior and user’s satisfaction level with the platform. Additionally, we include several scenarios related to “assuming interest but realizing a lack of it after further exploration” (i.e., tolerance, as we defined), to help respondents engage with “tolerance” scenarios in preparation for the next part of the survey. The questions are designed differently for e-commerce and content browsing applications.
Tolerance and Its Effect on Retention Rates. In this section, we ask questions about the impact of tolerance, specifically the accumulation of the phenomenon of “assuming interest but realizing a lack of it after further exploration”. A structured series of inquiries is used to evaluate how the gradual increase in tolerance toward recommended content affects the overall user experience on the platform, potentially leading users to permanently leave the platform.
@@ 3.4. Design of the Second User Survey
Table 1. List of 9 questions ($Q1$$Q9$) and their responses: Users can answer “yes” or “no” to these questions. Questions $Q1$$Q3$ are from the first survey, and the remaining questions are from the second survey.
@@
(a) Respondents of the first survey
(b) Respondents of the second survey
3.5. Survey Distribution and Collection
The user survey was distributed through CSDN [^1] https://www.csdn.net/, one of the largest developer communities in China. The first survey was launched on March 7, 2024, with the questionnaire link displayed in a banner on the CSDN homepage for 17 consecutive days. During this period, we also sent messages to all logged-in users, encouraging them to participate. Additionally, the survey was promoted through CSDN’s official account on WeChat, a widely used social media platform in China. A total of $6,644$ responses were collected. The second survey, conducted via the same platform and channels, ran for five days starting May 13, 2024, and received $3,864$ responses.
To ensure that the sample accurately represented the primary users of recommender systems, a preliminary screening question was included in both surveys. This question asked respondents whether they were aware of their engagement with recommender systems in their daily activities. We filtered out users who reported being unaware of recommender systems, resulting in $5,556$ and $2,879$ valid responses, for first and second survey respectively. We believe that the number of responses to both surveys is sufficient for reliable data analysis. Additionally, the basic demographic profiles of the users from both surveys are largely consistent, as reported in Figure 2.
4. Survey Analysis and Findings
We have collated and thoroughly analyzed the results of the two user surveys and drawn several conclusions. We found that clicking behavior does indicate a certain degree of user interest, but it ultimately depends on the user’s final judgment after obtaining sufficient information on the item, which is influenced by the platform’s interaction design. For example, on an autoplay video platform, users can explore content in depth without needing to click, and this is also a key way for users to express interest. Once users have gathered sufficient information about an item (e.g., after watching a video for a while), their final attitude becomes crucial. This is because the process of obtaining information involves a certain cost in terms of time or even money. If users invest this cost and are left with meaningless information, their goodwill and trust in the platform will diminish, leading to a decrease in user activity and retention.
In the following, we detail our findings from the survey results. Recall that the first three sections of questions are the same for both surveys, and we received very similar responses across both surveys for these questions, with the differences of ‘Yes’ rate smaller than 2%. If a question appears in both surveys, we report the numbers from the first survey due to the larger number of responses. However, the second survey obtained $2,879$ valid responses, which provides a sufficient sample size for drawing conclusions. Our findings are presented from four key perspectives: 1) the value proposition of recommender system, 2) click behavior and user attitudes, 3) presence of hesitation, and 4) the definition and impact of tolerance.
Note that, since our survey questions are designed for two scenarios, online shopping and content browsing (in the context of short-video sharing), the conclusions we draw apply to these two scenarios, and may or may not be generalizable to other contexts.
4.1. The value proposition of recommender system
@@
Based on the results summarized above, the primary requirement for recommender systems is to meet users’ need for relevant and valuable information during fragmented time slots, even when there is no explicit intention to acquire information.
4.2. Click behavior and user attitudes
@@ 4.3. Presence of hesitation
4.4. The definition and impact of tolerance
@@
Tolerant behaviors reduce user activity.:
From the survey results, 67% of users indicated that as the number of tolerable situations increased, their positive feelings toward the recommender system’s functionality decreased, leading them to question its effectiveness. Additionally, 70% of users reported that increased tolerance led to boredom, reduced interaction, and a likelihood of quitting the recommendation function, potentially resulting in platform abandonment.
Tolerant behaviors directly affect user retention.:
From the survey, 70% of users indicated that if tolerance continues to increase, they would leave the platform and not return. This leads us to an important conclusion of the study: the tolerance phenomenon from the user’s perspective directly impacts the platform’s retention rate.
Typical signals of tolerant behavior.:
According to the survey, 90% of users agree that in an online shopping scenario, clicking on a product without taking further actions, such as adding it to the cart, bookmarking, or placing an order, indicates a lack of genuine interest (see Table 2). Similarly, 88% of users agree that in video browsing, when a video is watched but not for a sufficient amount of time, or is viewed with actions like double-speed playback or skipping, it reflects a lack of interest (see Table 3). We interpret these behaviors as signs of diminished interest after initial engagement, adding to the user’s cognitive load. Consequently, we identify these signals as clear indicators of tolerance under the two surveyed scenarios.
@@
Attitude User interactions Percentage (%)
@@
Attitude User interactions Percentage (%)
@@
5. Tolerance Behavior
Based on our survey findings, reducing user tolerance and improving information acquisition efficiency can enhance retention rates in recommender systems. Thus, it is crucial to minimize recommendations that encourage tolerance behavior at a systemic level.
5.1. Identification of Tolerance Behavior
The user survey shows that tolerance signals begin with signs of interest and end with disinterest behaviors, as illustrated in Figure 4. Thus, treating a user’s interaction with a single item as a discrete session, the clear shift from interest to disinterest can be used to identify the tolerance process.
Although users generally exhibit interest in similar ways, specific methodologies are needed to quantify these behaviors, particularly for qualitative descriptors of “full engagement” with an item, which are challenging to measure. As discussed, initial item checking typically signifies the onset of interest, while engrossed browsing or other positive actions indicate continued interest. Conversely, ceasing to check suggests disinterest. The challenge lies in defining appropriate positive and negative signals. For example, in video recommendations, a browsing duration can be considered a positive signal of engagement, whereas durations reflecting “superficial engagement” (e.g., fast forwarding) are treated as negative indicators. However, individual differences, such as browsing habits and patience levels, add complexity. A highly patient user might engage with up to 80% of content they find uninteresting, while an impatient user might disengage after only 20%. This variability highlights the difficulty of applying uniform standards of engagement tolerance across different type of users. Additionally, the significance of engagement duration depends on the content’s total length. For instance, five seconds is substantial for a 10-second video but negligible for a 10-minute video.
Considering these factors, this study adopts the personalized framework of the recommender system, establishing individual-specific judgment criteria. We assume that the personalities of users, particularly long-term platform participants, are relatively stable and that their interaction patterns with content remain consistent. Therefore, in the context of video recommendations, it is reasonable to infer users’ attitudes from their average browsing ratios, a conclusion supported by Zannettou et al. (2024), which validated that users’ behavioral habits are stable. Additionally, in e-commerce, users’ attitudes can be inferred based on whether they favor or add items to their cart after clicking on them. Specifically, if a user does not engage with certain content at all, it is classified as a negative sample; if the user engages but the browsing ratio is below their average for similar content or shows no further positive behavior, it is considered superficial engagement and classified as a tolerance sample; if the user engages and the viewing ratio exceeds their average or there is further positive behavior, it is considered deep engagement and classified as a positive sample.
5.2. Offline Dataset Analysis: Tolerance Behaviors
According to the survey, over $60\%$ of participants reported that an increase in tolerance would lead to a decline in interest in the content. Consequently, they are likely to spent less time on the platform and may eventually discontinue its use. Here, we further understand user behaviors through a study conducted on two offline datasets. Specifically, we examined tolerance behaviors in Taobao [^2] https://tianchi.aliyun.com/dataset/42 and Kuaishou [^3] https://www.kuairand.com datasets, which represent two distinct platforms.
5.2.1. Datasets

### v2

@@ 2.3. User Retention Optimization
User retention is typically defined as the ratio of users who revisit a system after their initial interaction (Cai et al., 2023). It has become an increasingly critical metric in the evaluation and optimization of recommender systems (Wu et al., 2017; Zou et al., 2019; Cai et al., 2023; Xue et al., 2023b; Silveira et al., 2017; Zhang et al., 2021b; Ding et al., 2023; Gomez-Uribe and Hunt, 2016). Unlike immediate feedback, such as point-wise (He et al., 2021; Zamani and Croft, 2020) and list-wise (Cui et al., 2021; Zheng et al., 2020) signals, which reflect user responses after a single interaction, retention captures long-term feedback resulting from multiple interactions over time (Cai et al., 2023). This longer horizon aligns with prior findings that emphasize sustained user experiences and trust rather than single clicks or sessions.
@@
Due to the nature of delayed feedback, reinforcement learning (RL) methods are particularly well-suited to optimize user retention by addressing long-term rewards and user engagement (Zhang et al., 2022; Xue et al., 2023a; Zou et al., 2019). One of the representative works is RLUR (Cai et al., 2023), designed to address challenges such as uncertainty, bias, and delayed feedback, thereby improving retention by focusing on long-term positive outcomes. In addition, some studies have explored optimizing retention through RL by modeling mid- and long-term user behaviors (Wu et al., 2017; Zou et al., 2019; Xue et al., 2023b), investigating interpretable aspects of retention (Ding et al., 2023), or integrating exploratory behaviors with reinforcement learning to enhance user experiences (Chen et al., 2021). Recent efforts (Kohavi and Chen, 2024) also analyze the impact of false positives in online A/B tests for recommender systems, where the mined false positives can be regarded as retention signals.
While technical approaches frame retention as a long-term optimization objective, prior work has shown that returning to a system is not always a proxy for satisfaction: users may return out of habit, compulsion, or lack of alternatives rather than genuine engagement (Lukoff et al., 2021; DeVito et al., 2017). Thus, combining retention-oriented optimization with research on user agency, motivation, and trust highlights the importance of distinguishing between surface-level engagement and meaningful, value-aligned use. We focus on this synthesis by introducing hesitation and tolerance as behavioral signals that help bridge immediate and long-term perspectives on user-recommender interactions.
3. User Interaction States in RecSys
Hesitation and tolerance are inherently subjective states, involving psychological costs, emotional responses, and perceptions of wasted effort. In this paper, we define hesitation as a user state involving uncertainty and additional cognitive effort before deciding interest, and tolerance as a user state in which initial curiosity leads to unwanted engagement that users continue despite realizing disinterest. Such experiences are difficult to capture in behavioral logs, which record only observable actions.
Therefore, to address our first research question: whether an intermediate state exists between interest and disinterest in recommender systems, we conducted two large-scale online surveys. Surveys are well-suited to capture such psychological states, which are otherwise invisible in behavioral traces, and to establish their prevalence and subjective costs. Our goal was to examine everyday user experiences with recommender systems, focusing on whether intermediate interaction states can be identified, how users describe them, and what consequences they report. Rather than presenting participants with predefined labels such as hesitation or tolerance, we employed scenario-based questions to elicit their experiences in concrete forms (e.g., clicking without purchasing, speed-watching a video). The first survey established behavioral signals and validated our question framing, while the second replicated these measures with an independent sample and added items probing longer-term consequences such as frustration, reduced activity, and attrition. Together, the two surveys provide complementary evidence for understanding intermediate user-recommender interaction states.
We use tolerance in its dictionary sense of “enduring or putting up with something undesirable,” rather than the prosocial or interpersonal meaning of tolerance. We adopt this term because users frequently described experiences in which they continued engaging with content despite realizing it was not what they wanted, often accompanied by feelings of wasted time or frustration. Other similar alternatives include “unwanted engagement” and “regretful engagement,” and we chose “tolerance” to best capture the pattern of brief, low-value persistence observed in our data.
3.1. Method
We introduce how we conducted the two large-scale surveys. Both surveys were reviewed and approved by Alibaba’s internal research ethics and compliance process.
(a) Respondents of the first survey
(b) Respondents of the second survey
3.1.1. Participants
The surveys were distributed through CSDN [^1] Chinese Software Developer Network (CSDN, https://www.csdn.net), one of the largest developer communities in China, and promoted on both the CSDN website and its official WeChat account. The first survey ran for 17 days and received 6,644 responses. The second survey ran for five days and received 3,864 responses. We designed two surveys rather than one to balance exploration and validation. The first survey established behavioral signals and tested the validity of our constructs, while the second replicated these measures with an independent sample and extended them to examine longer-term consequences. Together, the two surveys provide both robustness and complementary insights.
To ensure relevance, both surveys included a screening item asking whether respondents were aware of recommender systems in their daily activities. Responses failing this screen were excluded, leaving 5,556 valid responses in the first survey and 2,879 valid responses in the second. Demographic information (age, education, occupation) was collected in each survey. The two samples were consistent in distribution, as shown in Figure 1. We acknowledge that this recruitment strategy favors technically oriented users, but the large sample sizes and consistency across two independent surveys provide a reliable foundation for our analysis. Demographic information such as age, education, and occupation was also collected, allowing us to assess representativeness and identify potential sampling biases.
3.1.2. Survey Design
We discuss how we framed hypotheses that guided our study, the principles of how we designed the survey, and the detailed context for both surveys.
Hypothesis
We framed four hypotheses to guide our study:
  • Hesitation as a distinct state. We expect that users often encounter an intermediary state, hesitation, before deciding whether they are interested. This state requires additional time and effort compared to immediate acceptance or rejection.
  • Costs of hesitation. We hypothesize that the effort invested during hesitation influences subsequent judgments of interest or disinterest, shaping the overall experience with recommended items.
  • From hesitation to tolerance. When hesitation ends in disinterest, we anticipate a secondary state of tolerance, marked by frustration over wasted effort. Compared to direct rejection, tolerance represents a more negative and costly user experience.
  • Accumulated impact of tolerance. Finally, we hypothesize that repeated tolerance erodes trust in the system, reduces engagement, and, over time, drives users toward attrition.
These hypotheses highlight the emotional complexity of interacting with recommender systems. By examining states such as hesitation and tolerance, we aim to surface subtle but consequential factors that shape how people experience recommendations and whether they continue to engage with a platform. The hypothesis part in Figure 2 illustrates this process. Recognizing and addressing these intermediate states opens opportunities to design systems that not only recommend effectively but also respect users’ time, reduce frustration, and ultimately foster greater trust and long-term retention.
Survey Design Principles
To examine these hypotheses, we designed two large-scale user surveys. Prior work on questionnaire design cautions that asking directly about abstract states, such as hesitation, often produces superficial or unreliable responses  (Krosnick, 2018; Schwarz, 1999). Following this guidance, we avoided using the terms hesitation or tolerance explicitly. Instead, we described realistic scenarios, such as “assuming interest but realizing otherwise after further exploration,” to elicit responses that reflect users’ lived experiences.
To reduce bias, we conducted two independent surveys at different times. Each was structured into five sections that gradually guided participants from general behaviors toward more reflective judgments. This sequential design encouraged participants to recall experiences step by step, improving both the reliability and validity of the responses.
In the first survey, we focused on capturing how people signal both interest and disinterest in recommendations, as well as the more ambiguous cases where something initially seems appealing but later proves uninteresting. To ground our question design, participants were asked to rate the extent to which they agreed with predefined options, helping us validate the reasonableness of our categories. We also probed whether respondents had experienced hesitation, and invited them to reflect on what triggered it. Finally, we presented scenarios in which initial curiosity gave way to disinterest after hesitation, asking participants to evaluate which of these mirrored their own experiences with recommender systems. The second survey replicated the first three sections to cross-check the robustness of our findings, but also introduced two new sections targeting the downstream effects of hesitation and tolerance. These questions examined whether hesitation shaped eventual judgments of interest, and whether repeated tolerance experiences reduced engagement or contributed to attrition.
The two surveys were administered independently at different times, with no intentional overlap in participants. Both included basic demographic items, i.e., age, occupation, and education level, to assess the representativeness of the sample and evaluate the health of the survey traffic. This demographic profiling helped us interpret the validity of the conclusions and ensured that patterns were not artifacts of a narrow subgroup.
First Survey
The first survey was organized into five sections, each aimed at capturing a different aspect of how people interact with recommender systems (see Appendix A.1).
User profile. We began by collecting demographic information such as age, education, and occupation. This helped us understand who our respondents were and evaluate the representativeness of the survey platform.
Context Introduction. To situate participants, we asked broad questions about their everyday experiences with recommender systems: what they appreciated, what they found frustrating, and how these systems fit into their routines. These reflections established a baseline of perceived value and shortcomings, which informed later sections.
User Behavior Survey. We then turned to concrete behaviors in two familiar domains: online shopping and short-video browsing. Participants described what they typically do when they are interested, disinterested, or when something initially seems appealing but later proves not to be. We also asked about incidental actions (e.g., curiosity clicks or trending topics) to examine how such behaviors might distort recommendation outcomes.
Description and Assessment of the “Hesitation” State. Next, we explicitly probed hesitation. Participants were asked whether they had encountered moments of uncertainty and what caused them, such as insufficient information, doubts about quality, or suspicion of advertising. This section encouraged respondents to recall and articulate hesitation as a psychological state in their experience.
Recognition of Behavioral Combinations. Finally, we presented hypothetical scenarios where initial interest gave way to disinterest in shopping and video contexts. By asking users to identify which scenarios matched their experiences, we aimed to validate tolerance behaviors as recognizable and relatable states, rather than abstract constructs.
@@ 3.1.2. Survey Design
Table 1. List of 9 questions (Q1Q1Q9Q9) and their responses: Users can answer “yes” or “no” to these questions. Questions Q1Q1Q3Q3 are from the first survey, and the remaining questions are from the second survey.
@@
Second Survey
The second survey was designed to complement and validate the first. The initial three sections repeated the same questions as the first survey, allowing us to check consistency across independent samples collected at different times. Beyond replication, two new sections were added to probe more deeply into the impact of hesitation and tolerance (see Appendix A.2).
Further Investigation into Hesitation. Here, we aim to test the fourth hypothesis in Section 3.1.2, we asked whether users tend to prefer items they have deliberated over or whether they trust their first impressions and avoid hesitation altogether. This allowed us to test our hypothesis that hesitation is inversely related to satisfaction. We also introduced scenarios of “assuming interest but realizing otherwise after further exploration,” preparing participants to reflect on tolerance as a recognizable experience. The scenarios were tailored to both e-commerce and short-video contexts.
Tolerance and Its Effect on Retention Rates. This section examined the cumulative consequences of tolerance. Participants were asked whether repeated experiences of wasted time made them less engaged, more likely to disengage, or even inclined to abandon the platform altogether. These questions helped us connect tolerance to longer-term outcomes such as trust, fatigue, and attrition.
3.1.3. Analysis
We analyzed survey responses using descriptive statistics for closed-ended questions and thematic analysis (Braun and Clarke, 2006) for open-ended responses. For binary questions, we calculated the proportion of “yes” responses to assess prevalence of each experience or behavior. To evaluate consistency across surveys, we compared agreement rates for overlapping questions (Q1-Q3) between the two independent samples. The first three sections were identical across both surveys, and we found highly consistent responses with differences in “yes” rates smaller than 2%. For questions appearing in both surveys, we report results from the first survey due to its larger sample size (5,556 vs. 2,879 valid responses), though both sample sizes provide sufficient statistical power for drawing conclusions. For open-ended responses describing hesitation triggers and tolerance experiences, we performed thematic analysis by reading through responses, identifying initial patterns, and developing a coding framework. This framework was then applied to categorize responses according to common themes related to our questions about intermediate interaction states.
3.2. Results and Findings
Our findings are organized around four key perspectives: (1) the value proposition of recommender systems, (2) click behavior and user attitudes, (3) presence of hesitation, and (4) the definition and impact of tolerance. We focus our analysis on two prevalent recommender system contexts: online shopping and content browsing in short-video platforms, as these represent common everyday interactions where users make rapid interest judgments.
3.2.1. The value proposition of recommender system
@@
Based on the results above, the primary requirement for recommender systems is to meet users’ need for relevant and valuable information during fragmented time slots, even when there is no explicit intention to acquire information.
3.2.2. Click behavior and user attitudes
@@ 3.2.3. Presence of hesitation
3.2.4. The definition and impact of tolerance
@@
Tolerant behaviors reduce user activity. From the survey results, 67% of users indicated that as the number of tolerable situations increased, their positive feelings toward the recommender system’s functionality decreased, leading them to question its effectiveness. Additionally, 70% of users reported that increased tolerance led to boredom, reduced interaction, and a likelihood of quitting the recommendation function, potentially resulting in platform abandonment.
Tolerant behaviors directly affect user retention. From the survey, 70% of users indicated that if tolerance continues to increase, they would leave the platform and not return. This leads us to an important conclusion of the study: the tolerance phenomenon from the user’s perspective directly impacts the platform’s retention rate.
Typical signals of tolerant behavior. According to the survey, 90% of users agree that in an online shopping scenario, clicking on a product without taking further actions, such as adding it to the cart, bookmarking, or placing an order, indicates a lack of genuine interest (see Table 2). Similarly, 88% of users agree that in video browsing, when a video is watched but not for a sufficient amount of time, or is viewed with actions like double-speed playback or skipping, it reflects a lack of interest (see Table 3). We interpret these behaviors as signs of diminished interest after initial engagement, adding to the user’s cognitive load. Consequently, we identify these signals as clear indicators of tolerance under the two surveyed scenarios.
@@
Attitude User interactions Pct. (%)
@@
Attitude User interactions Pct. (%)
@@
4. Analyzing Tolerance Behavior with Offline Datasets
While surveys reveal that hesitation and tolerance erode satisfaction and retention intentions, they remain subjective and subject to recall or reporting biases. Users may claim they would disengage, but whether such disengagement occurs in practice requires objective validation, motivating us to complement them with behavioral datasets. To address this limitation, we analyze large-scale behavioral logs from two contrasting domains, e-commerce (Taobao [^2] https://tianchi.aliyun.com/dataset/42) and short-video (Kuaishou [^3] https://www.kuairand.com). These datasets allow us to systematically observe tolerance behaviors, such as clicking without purchase or watching below one’s typical ratio, and examine their correlation with reduced engagement. These behavior signals are user-normalized and reflect systematic under-engagement relative to a user’s typical behavior; they therefore distinguish tolerance from random noise or low-quality items. Their consistency across users and their strong predictive value for reduced future engagement further support their interpretation as tolerance signals.
Through this behavioral analysis, we move beyond self-reported perceptions and address our second research question: how can this intermediate state be systematically identified and modeled?
4.1. Datasets